OpenStudy (anonymous):

A standard showerhead in Reyna's house dispenses 9 gallons of water per minute. Reyna changed this showerhead to an energy-saving one. The equation shows the amount of water dispensed, y, as a function of the number of minutes, x, for the new showerhead. y = 3x How much water does Reyna save each day with the change in showerheads if she uses the shower for 8 minutes each day? 14 gallons 48 gallons 69 gallons 96 gallons

OpenStudy (amistre64):

how many gallons are used in 8 minutes?

OpenStudy (anonymous):

@sleepyjess @Shadowgirl15 @Smita12 @aryandecoolest

OpenStudy (anonymous):

@sierra! @LegendarySadist @AbigaelRose

OpenStudy (anonymous):

savings= 8*9-8*3=72-24=48 gallons

OpenStudy (anonymous):

8*9=72 8*3=24 72-24=48 48 gallons :)